# An extension of Steinberg’s Theorem to biquotient pairs of subgroups

### v2

@@ 3 The proof
For the second equality, we use the fact that the dimension formula holds in $\mathrm{R}G$ (Lemma 11), and, a fortiori, in the localization $(\mathrm{R}G)_{\mathfrak{p}}$. The ideal $\mathfrak{p}^{\prime}$ is, in particular, contained in $\mathrm{Supp}_{\mathrm{R}G}\mathrm{R}H_{1}\cap\mathrm{Supp}_{\mathrm{R}G}\mathrm{R}H_{2}$, and hence by Proposition 22 (b) it is contained in $\mathrm{Supp}_{\mathrm{R}G}(H_{1}\cap gH_{2}g^{-1})$ for some $g\in G$. The final inequality therefore follows from Proposition 22 (c) and Lemma 6 (c).
On the other hand, $(\mathbb{x},i^{*}\mathbb{y})$ is a sequence of length
$c_{1}+c_{2}$ $=(\operatorname{rank}G-\operatorname{rank}H_{1})+(\operatorname{rank}G-\operatorname{rank}H_{2})$
=(${=}\operatorname{rank}G-\mathrm{interrank}(H_{1},H_{2}).$ ) .
As $(\mathrm{R}T)_{\mathfrak{p}}$ is regular (Lemma 15), and hence in particular Cohen-Macaulay, we can now apply [Hartshorne]*Theorem 8.21A (c) or [Matsumura:CA]*§ 16.B, Theorem 31 to deduce that the $(\mathbb{x},i^{*}\mathbb{y})$ is a regular sequence. Hence, using [Eisenbud]*Corollary 17.5, we conclude that the Koszul complex $\Lambda^{\bullet}_{(\mathrm{R}T)_{\mathfrak{p}}}((\mathrm{R}T)_{\mathfrak{p}};\mathbb{x},i^{*}\mathbb{y})$ is exact in all positive degrees. ∎
@@
Consider the triple $(G\times G,T_{1}\times T_{2},\Delta G)$. The group $G\times G$ is good as $G$ is good and $\mathrm{R}(G\times G)\cong\mathrm{R}(G)\otimes\mathrm{R}(G)$. The subgroups $T_{1}\times T_{2}$ and $\Delta G$ are connected. Using Proposition 25, we see that the triple still satisfies the maximal rank condition (2). The restriction $\Delta^{*}\colon\mathrm{R}(G\times G)\to\mathrm{R}(G)$ is surjective. Also, given that $G$ is good, we can easily find a regular sequence of length $\operatorname{rank}G$ generating the kernel of $\Delta^{*}$. So Proposition 32 applies to $(G\times G,T_{1}\times T_{2},\Delta G)$. Using Proposition 25, we conclude that $\operatorname{Tor}_{i}^{\mathrm{R}G}(\mathrm{R}T_{1},\mathrm{R}T_{2})_{\mathfrak{p}}=0$ for all $i>0$ and for all prime ideals $\mathfrak{p}\in\mathrm{Supp}_{\mathrm{R}G}\mathrm{R}(\mathrm{Z}G)$. ∎
@@
Choose maximal tori $T_{1}$ and $T_{2}$ of $H_{1}$ and $H_{2}$, respectively. Replacing $H_{1}$ and $H_{2}$ by conjugate subgroups if necessary, we may assume that $T_{1}$ and $T_{2}$ are contained in a joint maximal torus $T$ of $G$. Note that by Proposition 26, passing from the groups $H_{i}$ to their maximal tori $T_{i}$ does not change the intersection rank. By Proposition 29, we can find a torus $T_{1}^{+}$ such that $T_{1}\subseteq T_{1}^{+}\subseteq T$, such that $\mathrm{interrank}(T_{1}^{+},T_{2})=\mathrm{interrank}(T_{1},T_{2})$ and such that the triple $(G,T_{1}^{+},T_{2})$ satisfies the maximal rank condition (6). So by Proposition 33, $\operatorname{Tor}_{*}^{\mathrm{R}G}(\mathrm{R}(T_{1}^{+}),\mathrm{R}T_{2})_{\mathfrak{p}}$ vanishes in all positive degrees for all $\mathfrak{p}\in\mathrm{Supp}_{\mathrm{R}G}\mathrm{R}(\mathrm{Z}G)$
We can now apply Theorem 19 with $\Lambda$, $\Gamma$, $A$ and $C$ as follows:
@@
$\operatorname{Tor}_{i}^{\mathrm{R}G}(\mathrm{R}T_{2},\mathrm{R}T_{1})_{\mathfrak{p}}\cong\operatorname{Tor}_{i}^{\mathrm{R}T_{1}^{+}}(\dots,\mathrm{R}T_{1})_{\mathfrak{p}}$
in all degrees $i$. The module on the right vanishes for all $i>\operatorname{pd}_{\mathrm{R}T_{1}^{+}}\mathrm{R}T_{1}$, and by Corollary 18, $\operatorname{pd}_{\mathrm{R}T_{1}^{+}}\mathrm{R}T_{1}\leq\operatorname{rank}T_{1}^{+}-\operatorname{rank}T_{1}$. As $(G,T_{1}^{+},T_{2})$ satisfies (6), and as $\mathrm{interrank}(T_{1}^{+},T_{2})=\mathrm{interrank}(T_{1},T_{2})$, this bound can be rewritten as $\operatorname{rank}T_{1}^{+}-\operatorname{rank}T_{1}=\operatorname{rank}G-\operatorname{rank}T_{1}-\operatorname{rank}T_{2}+\mathrm{interrank}(T_{1},T_{2})$. We thus deduce that $\operatorname{Tor}_{i}^{\mathrm{R}G}(\mathrm{R}T_{1},\mathrm{R}T_{2})_{\mathfrak{p}}$ vanishes in the required range. Using Proposition 31, we obtain the same claim for the triple $(G,H_{1},H_{2})$. ∎
Theorem 3 is now just a special case:
Proof of Theorem 3.
Fix some degree $i>\operatorname{rank}G-(\operatorname{rank}H_{1}+\operatorname{rank}H_{2})+\mathrm{interrank}(H_{1},H_{2})$. Any prime $\mathfrak{p}$ supporting $\operatorname{Tor}_{i}^{\mathrm{R}G}(\mathrm{R}H_{1},\mathrm{R}H_{2})$ is necessarily contained in $\mathrm{Supp}_{\mathrm{R}G}\mathrm{R}H_{1}\cap\mathrm{Supp}_{\mathrm{R}G}\mathrm{R}H_{2}$. So by Proposition 22 (b), it is contained in $\mathrm{Supp}_{\mathrm{R}G}\mathrm{R}(H_{1}\cap gH_{2}g^{-1})$ for some $g\in G$. As $(G,H_{1},H_{2})$ satisfies the lax biquotient condition, each of the intersections $H_{1}\cap gH_{2}g^{-1}$ is contained in the centre $\mathrm{Z}G$. So by Proposition 22 (a) we have $\mathfrak{p}\in\mathrm{Supp}_{\mathrm{R}G}\mathrm{R}(\mathrm{Z}G)$, hence by Proposition 34, $\operatorname{Tor}_{i}^{\mathrm{R}G}(\mathrm{R}H_{1},\mathrm{R}H_{2})_{\mathfrak{p}}=0$. ∎
Finally, Theorem 1 is, of course, a special case of Theorem 3.
@@ 4 K-theory of biquotients
This short final section explains how to apply Theorem 1 to biquotients to obtain Corollary 2.
@@
$\{\operatorname{Tor}^{p}_{\mathrm{R}G}(\mathrm{K}_{G}^{*}(X),\mathrm{K}_{G}^{*}(Y))\}^{q}\Rightarrow\mathrm{K}^{p+q}_{G}(X\times Y).$
Here, the $\mathrm{R}G$-modules $\mathrm{K}_{G}^{*}(X)$ and $\mathrm{K}_{G}^{*}(Y)$ are understood to be $\mathbb{Z}/2$-graded, and $\operatorname{Tor}$ denotes graded Tor. The $G$-spaces are assumed to be compactly generated and have the homotopy type of a CW complex, see [Hodgkin, § I.1].
Given a strict biquotient $H_{1}\backslash G/H_{2}$, we have a principal $G$-bundle $H_{1}\backslash G\times G/H_{2}\to H_{1}\backslash G/H_{2}$ [Singhof:DCM]. Take $X:=H_{1}\backslash G$ and $Y:=G/H_{2}$. Then $\mathrm{K}^{*}_{G}(X)\cong\mathrm{R}H_{1}$, $\mathrm{K}^{*}_{G}(Y)\cong\mathrm{R}H_{2}$, and $\mathrm{K}^{*}_{G}(X\times Y)\cong\mathrm{K}^{*}(H_{1}\backslash G/H_{2})$. So the spectral sequence takes the form
$\operatorname{Tor}^{p}_{\mathrm{R}G}(\mathrm{R}H_{1},\mathrm{R}H_{2})\Rightarrow\mathrm{K}^{p}(H_{1}\backslash G/H_{2}).$
In particular, given that $\mathrm{R}H_{1}$ and $\mathrm{R}H_{2}$ are concentrated in degree zero, the graded Tor is just the usual Tor. Theorem 1 therefore implies the results on the K-theory of biquotients summarized in Corollary 2 in exactly the same way that Steinberg’s Theorem and more generally Corollary 18 provide a computation of the K-theory of homogeneous spaces. We refer to [SCSQ, proof of Theorem 3.6 and, in particular, Figure 1] for details of the argument.

### v3

@@ 3 The proof
$=c_{1}+c_{2}.$
For the second equality in this chain, we use the fact that the dimension formula holds in $\mathrm{R}G$ (Lemma˜12), and, a fortiori, in the localization [^4] For the localization $(\mathrm{R}G)_{\mathfrak{p}}$, the dimension formula may alternatively be deduced following the argument for $(\mathrm{R}T)_{\mathfrak{p}}$ presented in the next paragraph: $(\mathrm{R}G)_{\mathfrak{p}}$ is a regular local ring, hence Cohen–Macaulay, so a version of the dimension formula for arbitrary, not necessarily prime, ideals holds by [Matsumura:CRT, Theorem 17.4 (i)]. $(\mathrm{R}G)_{\mathfrak{p}}$, so that both sides of the equation can be identified with the dimension of $(\mathrm{R}G_{\mathfrak{p}})_{\mathfrak{p}^{\prime}}=\mathrm{R}G_{\mathfrak{p}^{\prime}}$. For the inequality, note that by (4) the ideal $\mathfrak{p}^{\prime}$ is contained in $\mathrm{Supp}_{\mathrm{R}G}(\mathrm{R}H_{1}\otimes_{\mathrm{R}G}\mathrm{R}H_{2})$, and that hence by Proposition˜23 (b) it is contained in $\mathrm{Supp}_{\mathrm{R}G}(H_{1}\cap gH_{2}g^{-1})$ for some $g\in G$. The inequality therefore follows from Proposition˜23 (c) and Lemma˜6 (c). The final equality is immediate from (2) and the definitions of $c_{i}$ as $\operatorname{rank}G-\operatorname{rank}H_{i}$. Note that $c_{1}+c_{2}$ is precisely the length of the sequence $(\mathbf{x},i^{*}\mathbf{y})$.
On the other hand, as $(\mathrm{R}T)_{\mathfrak{p}}$ is regular (Lemma˜16), and hence in particular Cohen–Macaulay [Matsumura:CRT, Theorem 17.8], we also know that the difference in dimensions on the left side of (7) is equal to the height of the ideal generated by $\mathbf{x}$ and $i^{*}\mathbf{y}$ [Matsumura:CRT, Theorem 17.4 (i)]. This height is bounded above by $c_{1}+c_{2}$, by Krull’s Hauptidealsatz. Thus, we obtain equality in (7), and conclude:
$\dim{(\mathrm{R}T)_{\mathfrak{p}}}-\dim{(\mathrm{R}T)_{\mathfrak{p}}/(\mathbf{x},i^{*}\mathbf{y})}=c_{1}+c_{2}$
Using once more that $(\mathrm{R}T)_{\mathfrak{p}}$ is Cohen–Macaulay, we can now apply [Hartshorne]*Chapter II, Theorem 8.21A (c) or [Matsumura:CRT, Theorem 17.4 (iii)] to deduce that the sequence $(\mathbf{x},i^{*}\mathbf{y})$ is regular. Hence, using [Matsumura:CRT, Theorem 16.5 (i)], we conclude that the Koszul complex $\Lambda^{\bullet}_{(\mathrm{R}T)_{\mathfrak{p}}}(\mathbf{x},i^{*}\mathbf{y})$ is exact in all positive degrees. ∎
@@
Consider the triple $(G\times G,T_{1}\times T_{2},\Delta G)$. The group $G\times G$ is good as $G$ is good and $\mathrm{R}(G\times G)\cong\mathrm{R}G\otimes\mathrm{R}G$. The subgroups $T_{1}\times T_{2}$ and $\Delta G$ are connected. Using Proposition˜27 (b), we see that the triple still satisfies the maximal rank condition (2). The restriction $\Delta^{*}\colon\mathrm{R}(G\times G)\to\mathrm{R}G$ is surjective. Also, given that $G$ is good, we can easily find a regular sequence of length $\operatorname{rank}G$ generating the kernel of $\Delta^{*}$. So Proposition˜35 applies to $(G\times G,T_{1}\times T_{2},\Delta G)$. Using Proposition˜27 (c) and (d), we conclude that $\operatorname{Tor}_{i}^{\mathrm{R}G}(\mathrm{R}T_{1},\mathrm{R}T_{2})_{\mathfrak{p}}=0$ for all $i>0$ and for all prime ideals $\mathfrak{p}\in\mathrm{Supp}_{\mathrm{R}G}\mathrm{R}(\mathrm{Z}G)$. ∎
@@
Choose maximal tori $T_{1}$ and $T_{2}$ of $H_{1}$ and $H_{2}$, respectively. By Proposition˜28, passing from the groups $H_{i}$ to their maximal tori $T_{i}$ does not change the intersection rank. By Proposition˜31, we can find a torus $T_{1}^{+}\supseteq T_{1}$ such that $\mathrm{interrank}(T_{1}^{+},T_{2})=\mathrm{interrank}(T_{1},T_{2})$ and such that the triple $(G,T_{1}^{+},T_{2})$ satisfies the maximal rank condition (8). So by Proposition˜36, $\operatorname{Tor}_{*}^{\mathrm{R}G}(\mathrm{R}(T_{1}^{+}),\mathrm{R}T_{2})_{\mathfrak{p}}$ vanishes in all positive degrees for all $\mathfrak{p}\in\mathrm{Supp}_{\mathrm{R}G}\mathrm{R}(\mathrm{Z}G)$.
We can now apply Theorem˜20 with $\Lambda$, $\Gamma$, $A$ and $C$ as follows:
@@
$\operatorname{Tor}_{i}^{\mathrm{R}G}(\mathrm{R}T_{2},\mathrm{R}T_{1})_{\mathfrak{p}}\cong\operatorname{Tor}_{i}^{\mathrm{R}T_{1}^{+}}(M,\mathrm{R}T_{1})_{\mathfrak{p}}$
in all degrees $i$, with $M:=\mathrm{R}T_{2}\otimes_{\mathrm{R}G}\mathrm{R}T_{1}^{+}$. The Tor groups on the right vanish for all $i>\operatorname{pd}_{\mathrm{R}T_{1}^{+}}\mathrm{R}T_{1}$, and by Corollary˜19, $\operatorname{pd}_{\mathrm{R}T_{1}^{+}}\mathrm{R}T_{1}\leq\operatorname{rank}T_{1}^{+}-\operatorname{rank}T_{1}$. As $(G,T_{1}^{+},T_{2})$ satisfies (8), and as $\mathrm{interrank}(T_{1}^{+},T_{2})=\mathrm{interrank}(T_{1},T_{2})$, this bound can be rewritten as $\operatorname{rank}T_{1}^{+}-\operatorname{rank}T_{1}=\operatorname{rank}G-\operatorname{rank}T_{1}-\operatorname{rank}T_{2}+\mathrm{interrank}(T_{1},T_{2})$. We thus deduce that $\operatorname{Tor}_{i}^{\mathrm{R}G}(\mathrm{R}T_{1},\mathrm{R}T_{2})_{\mathfrak{p}}$ vanishes in the required range. Using Proposition˜34, we obtain the same claim for the triple $(G,H_{1},H_{2})$. ∎
Theorem˜3 is now just a special case:
Proof of Theorem˜3.
Fix some degree $i>\operatorname{rank}G-(\operatorname{rank}H_{1}+\operatorname{rank}H_{2})+\mathrm{interrank}(H_{1},H_{2})$. Consider a prime $\mathfrak{p}$ supporting $\operatorname{Tor}_{i}^{\mathrm{R}G}(\mathrm{R}H_{1},\mathrm{R}H_{2})$. Any such prime necessarily lies in the support of both $\mathrm{R}H_{1}$ and $\mathrm{R}H_{2}$. As $(G,H_{1},H_{2})$ satisfies the lax biquotient condition, this implies that $\mathfrak{p}$ also lies in the support of $\mathrm{R}(\mathrm{Z}G)$, as noted in Example˜25. Hence by Proposition˜37, $\operatorname{Tor}_{i}^{\mathrm{R}G}(\mathrm{R}H_{1},\mathrm{R}H_{2})_{\mathfrak{p}}=0$. So we find that the localization of $\operatorname{Tor}_{i}^{\mathrm{R}G}(\mathrm{R}H_{1},\mathrm{R}H_{2})$ at any prime ideal of $\mathrm{R}G$ is zero, hence $\operatorname{Tor}_{i}^{\mathrm{R}G}(\mathrm{R}H_{1},\mathrm{R}H_{2})$ itself must vanish. ∎
Finally, Theorem˜1 is, of course, a special case of Theorem˜3.
@@ 4 K-theory of biquotients
This short final section explains how to apply Theorem˜1 to biquotients to obtain Corollary˜2.
@@
$\{\operatorname{Tor}_{-p}^{\mathrm{R}G}(\mathrm{K}_{G}^{*}(X),\mathrm{K}_{G}^{*}(Y))\}^{q}\Rightarrow\mathrm{K}^{p+q}_{G}(X\times Y).$
Here, the $\mathrm{R}G$-modules $\mathrm{K}_{G}^{*}(X)$ and $\mathrm{K}_{G}^{*}(Y)$ are understood to be $\mathbb{Z}/2$-graded, and $\operatorname{Tor}$ denotes graded $\operatorname{Tor}$. The $G$-spaces are assumed to be compactly generated and have the homotopy type of a CW complex; see [Hodgkin, § I.1].
[…]
In particular, given that $\mathrm{R}H_{1}$ and $\mathrm{R}H_{2}$ are concentrated in degree zero, the graded Tor is just the usual Tor. Theorem˜1 therefore implies the results on the K-theory of biquotients summarized in Corollary˜2 in exactly the same way that Steinberg’s Theorem and more generally Corollary˜19 provide a computation of the K-theory of homogeneous spaces. We refer to [SCSQ, proof of Theorem 3.6 and, in particular, Figure 1] for details of the argument.